# TacoGFN: Target-conditioned GFlowNet for Structure-based Drug Design

### v5

@@ 5.4 Ablation studies
Effects of using a larger dataset. Here, we study the effect of training the docking score predictor on a larger dataset for de novo drug design. As shown in Table 3, TacoGFN using the docking score predictor trained on the larger ZINCDock-15M dataset demonstrates significant improvement in average docking score and Novel Hit %. This confirms it is possible to leverage the easily generated large-scale docking score data to generate more novel and higher affinity molecules.
@@
Docking score DS ($\downarrow$) Top-10 DS ($\downarrow$) Novel
Model dataset Avg. Med. Avg. Med. Hit % ($\uparrow$) Hit % ($\uparrow$)
TacoGFN CrossDock-100k -8.25 -8.39 -9.97 -10.26 58.25 43.49
TacoGFN ZINCDock-15M -8.35 -8.53 -9.97 -10.22 60.55 56.29
Effects of pocket conditioning. To examine the effect of the proposed pocket conditioning for GFlowNet, we train a molecular generation policy unconditioned on pocket information. The docking score predictor is unchanged - meaning it still predicts a docking score for a molecule with pocket information. The results are shown in Table 4. We observe that the pocket-conditioned GFlowNet achieves higher docking scores compared to the GFlowNet without pocket conditioning. This ablation validates that our method is indeed leveraging pocket conditioning to learn a family of molecular distribution across different pocket structures.
Table 4: Effectiveness of pocket structure conditioning.
Docking score DS ($\downarrow$) Top-10 DS ($\downarrow$)
@@ 6 Conclusion
In this paper, we have investigated the problem of structure-based drug design. To address the limitations of methods based on distribution learning, we have framed pocket-conditioned molecule generation as a multi-objective RL task. We have proposed TacoGFN, a target-conditioned GFlowNet which explores the chemical space to generate novel molecules with high binding affinity and desired properties such as drug-likeness. We have further introduced a docking score predictor using pharmacophore priors, to quickly compute an affinity reward for molecules. Our experiments on the CrossDocked2020 benchmark have demonstrated that TacoGFN outperforms the state-of-the-art methods in terms of docking scores, percentage of hits, and percentage of novel hits. This demonstrates the potential of TacoGFN as a powerful tool for structure-based drug discovery. In future work, we plan to validate the top generated ligands for some clinically relevant protein pockets in-vitro, i.e. in wet-lab experiments.
Impact statements
This paper presents work whose goal is to advance Machine Learning methods for drug discovery. Such methods are increasingly being employed in the pharmaceutical industry, since they promise to greatly speed-up the lengthy process of drug discovery and to significantly reduce its huge cost. If that promise holds, these Machine Learning methods will benefit patients through better care and our society through a reduction of the economic burden of drug development.
@@ Appendix A Softwares
In this study, we used the open-sourced code for GFlowNet (Bengio et al., 2021), PharmacoNet (Seo & Kim, 2023) and GVP-GNN (Jing et al., 2021). Our models were implemented using the Pytorch (Paszke et al., 2019) and PyTorch Geometric (Fey & Lenssen, 2019) libraries, which enabled efficient training and evaluation. We utilized RDKit (Landrum et al., 2006), a widely-used chem-informatics library, to handle the molecular structures and compute chemical properties. We employed the QuickVina 2.1 (QVina) (Alhossary et al., 2015) for docking, and used Openbabel (O’Boyle et al., 2011) and AutoDock Tools (Huey et al., 2012) to generate ready-to-dock files.
Appendix B Additional details of pocket conditioned GFlowNet
B.1 Action sampling
We present the action sampling process below:
$\mathbf{a}_{Add}$ $=MLP\left(\mathbf{h}^{L(N)}\right)$ (6)
$\mathbf{a}_{Attach}$ $=MLP\left(\mathbf{e}^{L(N)}\right)$ (7)
$\mathbf{a}_{Stop}$ $=MLP\left(\mathbf{g}^{L(N)}\right)$ (8)
@@ B.1 Action sampling
$\pi(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ $=softmax\left(Concat\left(\mathbf{a}_{Add},\mathbf{a}_{Attach},\mathbf{a}_{Step}\right)\right)$ (9)
$a$ $\sim\pi(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ (10)
B.2 Reward function
We normalize SA score using the formula $\frac{10-SA}{9}$ to obtain $\widehat{SA}$. We also multiply the predicted docking score by $-1$ to obtain a positive affinity reward $\widehat{DS}$. While $QED$ and $SA$ need to meet a certain threshold to make a good drug molecule, optimizing these values beyond the threshold does not bring additional utility (Coley, 2021). Therefore, we give no additional rewards to a molecule for having a $QED$ or $SA$ beyond $QED_{sat}$ and $SA_{sat}$ respectively. As the docking scores that indicate high affinity differ across pockets, we first compute $\overline{DS}$, which is the mean docking score of 1,000 random ZINC molecules for the given pocket, also multiplied by $-1$. Next, we normalize the docking score improvement of a given molecule and pocket by subtracting $\overline{DS}$ from $\widehat{DS}$. We clip this docking score improvement compared to a random molecule to be $\geq 0$. To avoid a constant 0 affinity improvement reward for all molecules that perform below average, we introduce a hyperparameter $\gamma$, which scales the docking score reward gained from docking score without mean subtracted. We use the following settings for all experiments: $QED_{sat}=0.7$, $SA_{sat}=0.8$, and $\gamma=0.2$. Formally the reward is defined as follows:
$QED_{reward}$ =$=\begin{dcases}1,&\text{if }QED>QED_{sat}\\ \frac{QED}{QED_{sat}},&\text{else}\end{dcases}$
$SA_{reward}$ =$=\begin{dcases}1,&\text{if }SA>SA_{sat}\\ \frac{SA}{SA_{sat}},&\text{else}\end{dcases}$
$DS_{reward}$ =$=\begin{dcases}\widehat{DS}-\overline{DS}+\gamma\widehat{DS},&\text{if }\widehat{DS}<\overline{DS}\\ \gamma\widehat{DS},&\text{else}\end{dcases}$
$Reward$ $=QED_{reward}\times SA_{reward}\times DS_{reward}$
B.3 Model Details
@@
Temp sampling distribution $Uniform(0,64)$
@@
Appendix C Additional details of docking score predictor
C.1 Motivation
@@
C.2 Training details
@@
C.3 Performance of docking score prediction model
Test set Training set RMSE (kcal/mol) MAE (kcal/mol) Pearson Spearman
CrossDocked-100k-test CrossDocked-100k 0.881 0.652 0.918 0.913
ZINCDock-15M 1.143 0.880 0.867 0.828
ZINCDock-test CrossDocked-100k 1.121 0.791 0.749 0.790
ZINCDock-15M 0.862 0.574 0.805 0.846
@@ D.1 Dataset
D.2 Docking protocol
In some previous evaluation protocols, the generated 3D molecules from the model are directly used as the conformations for the re-docking process. However, these generated 3D conformations may contain unrealistic bond length or bond angles (Harris et al., 2023) and therefore produce distorted docking scores. It is more reasonable and realistic to first calculate a conformer with realistic geometry, which is a necessary step in standard docking protocols, before docking (Bender et al., 2021). For our protocol, we first convert all generated molecules into SMILES and calculate their ETKDG conformers (srETKDGv3) using RDKit. Then, we prepare ready-to-dock files of ligands and proteins with Openbabel and AutoDockTools (O’Boyle et al., 2011; Huey et al., 2012). Finally, we dock these conformers using Quick VINA 2.1 (QVina) (Trott & Olson, 2010; Alhossary et al., 2015). For QVina, we use a box size of 20 Å and an exhaustiveness of 8.
D.3 Experimental metric details
Some baseline methods did not generate at least 10 molecules which satisfy the desirable criteria for a number of pockets. To compute Top-10 Desr. DS in these cases, a docking score of 0 is used as a placeholder for the top 10 docking scores.
D.4 Baseline evaluations
@@
D.5 Additional Results
D.5.1 Generated samples
Figure 5: To present of fair a overview of the model performance, we selected 4iwq and 4q8b at the $25$-th and $75$-th percentiles based on their docking scores with native ligands. We show compare the molecules generated by TacoGFN against the native ligand with their QED, SA, Novelty, and Docking score (Vina).
D.5.2 Generation quality
To measure the quality of molecule generation, we generated 100 molecules for each test pocket with state-of-the-art methods. Note that we do not resample the molecules until the number of valid and novel molecules reaches 100, in contrast to Section 5.2. Validity is defined as being readable by RDKit and successful in docking. We measured the average percentages of unique and valid molecules in the generated samples.
Table 8: The valid & unique is the average percentage of valid and unique molecules out of the 100 generated samples.
Model valid & unique (%)
Pocket2Mol (Peng et al., 2022) 71.86
TargetDiff (Guan et al., 2023a) 91.45
DecompDiff (Guan et al., 2023b) 66.32
TacoGFN 100.00

### v6

@@ 5.5 Ablation studies
We conduct additional ablation studies using the base TacoGFN under the generative setting to examine the effect of docking score prediction accuracy and pocket conditioning on our method.
Effects of using higher quality docking score predictor. Here, we study the effect of using a more accurate docking score predictor, which is trained on a larger dataset, for reward. Since training a docking score predictor does not require high-quality protein-ligand structural data such as the CrossDock-100k set, we can introduce a second, larger dataset for docking score prediction called ZINCDock-15M. It consists of about 15M docking simulation data - from docking 1,000 random ZINC20 (Irwin et al., 2020) molecules into each of the 15,207 unique pockets from CrossDock-100k training split using QVina. We then train our pharmacophore-based docking score predictor on this larger dataset. Please see Appendix Table 11 for a comparison of docking score accuracy between using CrossDock-100k and ZINCDock-15M. As shown in Table 5, TacoGFN using the docking score predictor trained on the larger ZINCDock-15M dataset demonstrates improvements in average Vina Dock, high-affinity rate and success rate. This confirms it is possible to leverage the easily generated large-scale docking score data to generate more novel and higher affinity molecules.
@@
Docking score Vina Dock ($\downarrow$) High Affinity($\uparrow$) Success Rate($\uparrow$)
Model dataset Avg. Med. Avg. Med. Avg. Med.
TacoGFN CrossDock-100k -8.24 -8.44 67.5% 92.0% 56.0% 61.5%
TacoGFN ZINCDock-15M -8.35 -8.53 69.5% 94.5% 58.3% 67.5%
Effects of pocket conditioning. To examine the effect of the proposed pocket conditioning for GFlowNet, we train a molecular generation policy unconditioned on pocket information. The docking score predictor is unchanged - meaning it still predicts a docking score for a molecule with pocket information. The results are shown in Table 6. We observe that the pocket-conditioned GFlowNet achieves higher docking scores compared to the GFlowNet without pocket conditioning. We further measure the number of non-covalent interactions of a molecule to respective pocket, by obtaining the binding pose using QVina (See Appendix D.4.1). We demonstrate generated molecules by pocket-conditioned GFlowNet result in more non-covalent interactions of for all categories (Hydrophobic interactions, Van der Waals contacts, Hydrogen binding) in Table 9. More non-covalent interactions are indicative of the molecule having better specificity to the protein target. This ablation validates that our method is indeed leveraging pocket conditioning to learn a family of molecular distribution across different pocket structures.
Table 6: Effectiveness of pocket structure conditioning evaluated using Vina Dock.
Docking score Vina Dock ($\downarrow$) Top-10 Vina Dock ($\downarrow$)
@@ 6 Conclusion
In this paper, we have investigated the problem of structure-based drug design. To address the limitations of methods based on distribution learning, we have framed pocket-conditioned molecule generation as learning a multi-objective reward distribution using RL. To this end, we propose TacoGFN, a pocket structure conditioned GFlowNet which generate drug-like and high-affinity molecules with respect to any 3D pocket structure. To the best of our knowledge, TacoGFN is the first RL model to address the challenging task of modelling a family of reward functions induced from all pocket structures. Our model effectively explores the greater chemical space, through generating millions of molecules using the online policy during training. We have further introduced a novel pharmacophore-based affinity predictor, where coarse-graining to the protein pocket is shown to achieve more accurate and robust predictions than existing architectures and protein representations. We finally introduce TacoGFN+FT, which fine-tunes the generic TacoGFN for a given test pocket.
Our experiments on the CrossDocked2020 benchmark have demonstrated that TacoGFN and TacoGFN+FT outperform the state-of-the-art methods in terms of Vina Dock, high affinity and success rate. This demonstrates the potential of TacoGFN as a powerful tool for structure-based drug discovery. In future work, we plan to validate the top generated ligands for some clinically relevant protein pockets in-vitro, i.e. in wet-lab experiments.
Broader Impact Statement
This paper presents work whose goal is to advance machine learning methods for drug discovery. Such methods are increasingly being employed in the pharmaceutical industry since they promise to greatly speed-up the lengthy process of drug discovery and to significantly reduce its huge cost. If that promise holds, these machine-learning methods will benefit patients through better care and our society through a reduction of the economic burden of drug development.
@@ Appendix A Softwares
In this study, we used the open-sourced code for GFlowNet (Bengio et al., 2021), PharmacoNet (Seo & Kim, 2023) and GVP-GNN (Jing et al., 2021). Our models were implemented using the Pytorch (Paszke et al., 2019) and PyTorch Geometric (Fey & Lenssen, 2019) libraries, which enabled efficient training and evaluation. We utilized RDKit (Landrum et al., 2006), a widely-used chem-informatics library, to handle the molecular structures and compute chemical properties. We employed the QuickVina 2.1 (QVina) (Alhossary et al., 2015) and UniDock (Yu et al., 2023) for docking, and used Openbabel (O’Boyle et al., 2011) and AutoDock Tools (Huey et al., 2012) to generate ready-to-dock files.
Appendix B Problem definition compared to existing RL baselines
TacoGFN addresses the problem of structure-based drug design (SBDD), which aims to have one model generate high-affinity molecules conditioned on any unseen protein structure. We train this one model to generate molecules conditional to different protein target structures in the CrossDocked training set, using predicted affinity, Synthetic Accessibility (SA), and drug-likeness (QED) as the reward. Then we evaluate the trained model’s performance on the 100 unseen protein target structures in the CrossDocked test set. Molecules are generated for unseen protein structures during evaluation to measure whether SBDD models have learned generalized protein-ligand interaction patterns during training.
On the other hand, many existing optimization-based RL methods (Bengio et al., 2021; Zhavoronkov et al., 2019; Korshunova et al., 2022; Jeon & Kim, 2020; Goel et al., 2021a; Lee et al., 2023; Reidenbach, 2024) focus on the target-free problem setting - meaning the target protein structure is not used as input conditioning. RGA Fu et al. (2022), DecompOpt Zhou et al. (2024) and TacoGFN+FT use protein structure as input, however, they still require docking oracle calls to the test protein pocket during evaluation to generate a molecule set. Therefore, they also fall within the optimization-based problem setting. These models are trained or optimized to generate molecules which optimize for predicted affinity on one protein target only and then evaluate on the same target. The goal of this evaluation is to measure the effectiveness of the algorithm in terms of generating objects which match a reward distribution. Although both methods use the commonly used metrics such as affinity, Synthetic Accessibility (SA), and drug-likeness (QED) as the reward, the goal of the evaluation and the problem difficulty is different compared to the SBDD setting.
Because protein-ligand interaction is highly specific, the distribution of molecules with high affinity will vary greatly across different possible pocket structures. The SBDD setting, which requires modelling the molecule distribution conditional to any protein pocket and generalizing to unseen protein pockets, is therefore a different and more challenging task than modelling molecule distribution for a single protein pocket.
Appendix C Method Details
C.1 Additional details of pocket conditioned GFlowNet
C.1.1 Molecular generation actions
We follow action used in previous works on fragment-based molecular generation (Jin et al., 2018; Bengio et al., 2021; Hamidizadeh et al., 2023) to construct molecules fragment by fragment. We present the three types of actions available to TacoGFN below:
(1) FragmentAddition: At each step, for each fragment node $v^{L}_{i}$ in the molecular graph, we apply the same MLP over its node embedding $\mathbf{h}^{L(N)}_{i}$ which produces logits over the fragment vocabulary. Each logit represents the unnormalized score for attaching fragment node $v^{L}_{i}$ to a particular new fragment node ($v^{L}_{j}$) from the vocabulary. These logits correspond to the FragmentAddition action type. While the FragmentAddition specifies whether fragment node $v^{L}_{i}$ connects to fragment node $v^{L}_{j}$, it does not specify how they are connected (i.e. which atom on fragment $v^{L}_{i}$ forms a bond with which atom on fragment $v^{L}_{j}$).
$\mathbf{a}_{Add}$ $=MLP\left(\mathbf{h}^{L(N)}\right)$
(2) AttachmentSpecification determines how the fragment pairs are connected. At each step, for each directional edge in the molecular graph which connects fragment $v^{L}_{i}$ to $v^{L}_{j}$, we produce a logit over the atoms of fragment $v^{L}_{i}$. Each logit represents the unnormalized score for fragment i connecting to fragment $v^{L}_{j}$ via a single bond from a particular atom on fragment $v^{L}_{i}$, based on edge embedding $\mathbf{e}^{L(N)}_{ij}$. The molecule can only be completed when all attachment edges are specified.
$\mathbf{a}_{Attach}$ $=MLP\left(\mathbf{e}^{L(N)}\right)$
(3) StopConstruction is a graph action that marks the finish of a molecule. The logit is produced from a single MLP output based on the final graph embedding $\mathbf{g}^{L(N)}$. All logits are concatenated and scaled into probabilities using the softmax function, and an action is sampled from the distribution. The same process is repeated for each time step until the stop construction action is sampled.
$\mathbf{a}_{Stop}$ $=MLP\left(\mathbf{g}^{L(N)}\right)$
After all the scores for all possible actions are computed, a final action is sampled based on the equations as follows:
@@ C.1.1 Molecular generation actions
$P_{a}(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$ $=softmax\left(Concat\left(\mathbf{a}_{Add},\mathbf{a}_{Attach},\mathbf{a}_{Step}\right)\right)$
$a$ $\sim P_{a}(\mathbf{a}|\mathcal{G}^{L}_{t},\mathbf{h}_{\mathcal{G}^{\mathcal{P}}})$
C.1.2 Reward function.
Our reward function consists of properties highly relevant for a drug candidate: Vina Docking Score (DS), Drug Likeliness (QED) and Synthetic Accessibility (SA). Unless otherwise mentioned, $t_{DS}=-8.0$, $t_{QED}=0.7$ and $t_{SA}=0.8$ for all models.
Drug Likeliness (QED) and Synthetic Accessibility (SA):
We first normalize the raw SA score using the formula $\frac{10-SA}{9}$ to obtain a reward between 0 and 1. While QED and SA need to meet a certain threshold to make a good drug molecule, optimizing these values beyond the threshold does not bring additional utility (Coley, 2021). Therefore, we clip reward $r$ for QED or SA to 1 when they achieve their respective threshold $t$. In other words, our model will not be incentivised to optimize QED/SA beyond their threshold value; Therefore more priority will be placed on optimizing the Vina Docking Score when these thresholds are reached.
$r_{QED}$ $=\min\left(\frac{QED}{t_{QED}},1\right)$
$r_{SA}$ $=\min\left(\frac{SA}{t_{SA}},1\right)$
Vina Docking Score (DS):
[…]
$r_{DS}$ $=-\frac{(DS-t_{DS})+0.2\times\max\left(DS,t_{DS}\right)}{\sqrt[3]{HAC}}$
We obtain the final reward by multiplying the normalized rewards together:
$r=r_{DS}\times r_{QED}\times r_{SA}$
Model Details.
@@
Training reward temp $\beta$ $Uniform(0,64)$
Inference reward temp 64
@@
C.2 Additional details of docking score predictor
Motivation.
@@ C.2 Additional details of docking score predictor
In the table, we find TacoGFN (C-alpha) outperforms BigBind and DeepBindGCN, indicating that our model architecture (see equation 5 of our paper) which integrates both pairwise interactions and global pooling information is more effective than the other architectures. Moreover, we also observed that the pharmacophore representation significantly improved the performance over the residue graph representation. Notably, TacoGFN (Pharmacophore) trained with CrossDocked-100k shows better performance on ZINCDock-test compared to the other models trained with ZINCDock-15M. Given that the main challenge of non-structure-based bioactivity predictors is poor generalization performance to ligands not used in the training set Chan et al. (2023), this result demonstrates that the pharmacophore representation of the pocket greatly improves both accuracy and generalization compared to the residue graph representation.
Training details
@@ D.1 Dataset
D.2 Docking protocols
For our protocol under the generative setting, we first convert all generated molecules into SMILES and calculate their ETKDG conformers (srETKDGv3) using RDKit. Then, we prepare ready-to-dock files of ligands and proteins with Openbabel and AutoDockTools (O’Boyle et al., 2011; Huey et al., 2012). Finally, we dock these conformers using Quick VINA 2.1 (QVina) (Trott & Olson, 2010; Alhossary et al., 2015). For QVina, we use a box size of 20 Å and an exhaustiveness of 8. Note for the optimization setting, we adopt UniDock instead following previous works (Reidenbach, 2024) (Please see details in Appendix E.1).
D.3 Baseline evaluations
@@
D.4 Additional Results
D.4.1 Additional ablation on effects of pocket conditioning
To further demonstrate the effectiveness of pocket conditioning, we compare the numbers of non-covalent interactions achieved by the generated molecules from pocket-conditioned TacoGFN with TacoGFN without pocket conditioning. The non-covalent interactions between protein and ligand fall into the following categories: (1) Hydrophobic Interactions are interactions between nonpolar regions of the protein and the ligand. (2) Van der Waals Contacts are weak forces arising from induced electrical interactions between closely positioned atoms of the protein and ligand. (3) Hydrogen Bonding refers to strong dipole-dipole interactions between an electronegative atom and (typically) a hydrogen atom. (4) Total Interactions refers to the total number of interactions between the protein and the ligand, including all types of non-covalent interactions. The generated molecules are docked using QVina (Trott & Olson, 2010; Alhossary et al., 2015) to generate a binding conformation. We use PoseCheck (Harris et al., 2023) to compute the number of interactions.
In table 9, we found that pocket conditioning increased all types of non-covalent interactions between protein and generated molecules. this result suggests that the pocket structural information can provide guidance in selecting functional groups that can form better supramolecular interactions with the target pocket.
Table 9: Effectiveness of pocket structure conditioning measured using the average numbers of non-covalent interactions (NCI) achieved after docking the generated molecules into the target pocket using QVina. Higher numbers of interactions indicate a stronger binding interaction.
Model Docking score dataset Hydrophobic ($\uparrow$) Van der Waals contacts ($\uparrow$) Hydrogen bonding ($\uparrow$) Total interactions ($\uparrow$)
TacoGFN ZINCDock-15M 5.67 7.23 1.15 14.05
w/o pocket conditioning ZINCDock-15M 5.19 6.83 1.11 13.14
D.4.2 Discussion on diversity
In our experiments with GFlowNet, we find a trade-off between diversity and the goal of optimizing for a higher average affinity of the generated candidate set. As TacoGFN seeks to generate molecules with a strong affinity to a specific pocket structure while optimizing for QED and SA, the resulting solution set is inevitably smaller, leading to reduced diversity.
We note that the molecular diversity of TacoGFN, if desired, can be increased in various ways. Firstly, the reward is currently exponentiated by reward temperture $\beta$. As $\beta$ increases, the reward density is concentrated close to the modes with high reward, and subsequently, there’s a reduction in diversity (Jain et al., 2023). Diversity can be improved by lowering the $\beta$.
Table 10: Comparison of average Vina Dock and diversity of TacoGFN trained under different reward settings.
Model Vina Dock ($\downarrow$) Diversity ($\uparrow$)
TacoGFN -8.24 0.53
$\textsc{TacoGFN}_{lenient}$7.91 0.73
In addition, we have experimented with various reward function settings. In the reward scenario adopted in the paper - the strict setting (see Appendix C.1.2), we reduce the reward given to the component docking score that is below -8.0 kcal/mol. We have found that the stricter reward produces better affinity metrics at the expense of lower diversity. In the lenient setting, where we simply reward docking scores based on their normalized raw values, $\textsc{TacoGFN}_{lenient}$ still outperforms Pocket2Mol and TargetDiff in terms of average docking score (-7.91) while having higher diversity (0.73).
In summary, if the key concern is candidate diversity and novelty, TacoGFN can be trained to generate generally high reward molecules under the lenient setting and/or low $\beta$. If the goal is to generate the best compound set with high average affinity and top affinity with decent diversity, and satisfying QED and SA, TacoGFN can be rewarded under the strict setting and/or high $\beta$.
Figure 5: The average of the top-10 Vina Dock of molecules generated for individual CrossDocked test pockets (target) by DecompDiff and TacoGFN. Color is used to denote the average molecular weight of molecules in the Top-10 set. Molecular mass of an orally active drug should be less than 500 daltons Lipinski et al. (1997); Heavy molecules with high docking scores are more likely to be false positives (Pan et al., 2002). Overall TacoGFN consistently achieves ideal molecular weight and strong Vina Dock.
D.4.3 Performance of docking score prediction model
While many existing affinity-prediction architectures either work for one target only (Bengio et al., 2021; Gentile et al., 2020), or require binding complex structures (Shen et al., 2022), we cannot adopt these model architectures as we require affinity prediction across various protein targets without using their binding complex structure. Thus, we compare our pharmacophore-based affinity prediction architecture with BigBind (Brocidiacono et al., 2022) and DeepBindGCN (Zhang et al., 2023a), two recently proposed non-complex-based methods which achieved state-of-the-art performance in virtual screening and affinity prediction, respectively. In addition, we examine the effect of introducing pharmacophore-based coarse-grained encoding of the protein pocket, by creating an ablation where we use residue graphs with C-alpha as nodes (as typically done in other methods) Lu et al. (2022) instead of pharmacophore graphs to represent our protein pocket. We call this ablation model TacoGFN (C-alpha). We re-train these models using the same CrossDocked2020 and ZINCDock-15M datasets and evaluate them on both test sets.
Model Training set CrossDocked-100k-test ZincDock-15M-test
RMSE MAE RMSE MAE
BigBind CrossDocked-100k 1.252 0.991 1.516 1.175
ZINCDock-15M 1.561 1.227 1.233 0.955
DeepBindGCN CrossDocked-100k 1.739 1.366 1.912 1.531
ZINCDock-15M 1.877 1.470 1.409 1.066
TacoGFN (C-alpha) CrossDocked-100k 1.091 0.821 1.517 1.157
ZINCDock-15M 1.483 1.156 1.170 0.872
TacoGFN (Pharmacophore) CrossDocked-100k 0.881 0.652 1.121 0.791
ZINCDock-15M 1.143 0.880 0.862 0.574
Appendix E Finetuning of TacoGFN
Table 12: TacoGFN+FT ablations: Here, we vary the reward function parameters.
Validity ($\uparrow$) Vina Dock ($\downarrow$) High Affinity ($\uparrow$) QED ($\uparrow$) SA ($\uparrow$) Diversity ($\uparrow$) Success Rate ($\uparrow$) Time ($\downarrow$)
Model Avg. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Gen + Score.
Reference 100% -7.45 -7.26 - - 0.48 0.47 0.73 0.74 - - 25.0% -
RGA - -8.01 -8.17 64.4% 89.3% 0.57 0.57 0.71 0.73 0.41 0.41 46.2% -
TargetDiff+Opt - -8.30 -8.15 71.5% 95.9% 0.66 0.68 0.68 0.67 0.31 0.30 25.8% >3728
DecompOpt - -8.98 -9.01 73.5% 93.3% 0.48 0.45 0.65 0.65 0.60 0.61 52.5% 9241
EvoSBDD ($\alpha$ = 1.3, 140R) 100% -10.27 -10.36 96.5% 100% 0.53 0.52 0.75 0.77 0.63 0.63 78.8% 6300
EvoSBDD ($\alpha$ = 0, $\sigma$ = 1, 140R) 100% -10.14 -10.27 94.4% 100% 0.59 0.59 0.77 0.77 0.62 0.62 86.4% 6300
TacoGFN+FT ($t_{QED}=0.40,t_{SA}=0.60,n=240$) 100% -10.63 -10.78 97.1% 100% 0.48 0.47 0.72 0.71 0.62 0.62 86.5% 6200
TacoGFN+FT ($t_{QED}=0.50,t_{SA}=0.80,n=240$) 100% -10.24 -10.31 97.0% 100% 0.58 0.58 0.82 0.82 0.61 0.61 87.7% 4980
TacoGFN+FT ($t_{QED}=0.50,t_{SA}=0.75,n=240$) 100% -10.21 -10.31 96.2% 100% 0.58 0.57 0.80 0.81 0.62 0.61 86.9% 5210
TacoGFN+FT ($t_{QED}=0.55,t_{SA}=0.80,n=240$) 100% -10.16 -10.27 96.8% 100% 0.62 0.61 0.82 0.82 0.60 0.60 86.8% 5010
TacoGFN+FT ($t_{QED}=0.60,t_{SA}=0.80,n=240$) 100% -10.07 -10.11 95.8% 100% 0.65 0.64 0.82 0.82 0.60 0.60 85.7% 4910
TacoGFN+FT ($t_{QED}=0.40,t_{SA}=0.60,n=300$) 100% -10.78 -10.93 97.1% 100% 0.47 0.47 0.70 0.69 0.62 0.62 87.8% 7750
TacoGFN+FT ($t_{QED}=0.50,t_{SA}=0.80,n=300$) 100% -10.32 -10.39 97.0% 100% 0.58 0.58 0.82 0.82 0.61 0.61 88.8% 6230
TacoGFN+FT ($t_{QED}=0.50,t_{SA}=0.75,n=300$) 100% -10.31 -10.41 96.2% 100% 0.57 0.57 0.80 0.80 0.62 0.62 87.7% 6520
TacoGFN+FT ($t_{QED}=0.55,t_{SA}=0.80,n=300$) 100% -10.24 -10.35 97.0% 100% 0.62 0.61 0.83 0.82 0.60 0.60 87.3% 6270
TacoGFN+FT ($t_{QED}=0.60,t_{SA}=0.80,n=300$) 100% -10.14 -10.21 96.2% 100% 0.65 0.65 0.82 0.82 0.59 0.59 86.3% 6140
E.1 Finetuning settings.
For finetuning TacoGFN, we adopt the same reward function as described in appendix section C.1.2. However, instead of using the previously mentioned docking score predictor for affinity reward, we use the docking program directly. Following the same setting as EvoSBDD Reidenbach (2024), we adopt UniDock (Yu et al., 2023) - a GPU-accelerated docking program to compute docking score between the generated molecule and the protein target during the fine-tuning stage. We use UniDock’s default balanced mode, which has an exhaustiveness of 384 and a max step of 40. We fine-tune using a batch size of 64 to leverage the speedup from UniDock’s parallelization for computing docking scores.
For a fair comparison with the previous methods, we re-scored the UniDock’s docking pose with Vina (in a score-only mode without any structural changes) and obtained the same values within the margin of error ($\pm$ 0.005 kcal/mol). The top 100 molecules with the highest rewards from the fine-tuning stage are used for evaluation.
E.2 Results.
For ablations studies, we vary our reward function by adjusting $t_{QED}$ and $t_{SA}$. Once the threshold for $t_{QED}$ or $t_{SA}$ is reached for molecules, the model will not be incentivised to optimize its QED or SA further. With a lower threshold standard ($t_{QED}$ and $t_{SA}$), our model will prioritize optimizing for Vina Dock instead of optimizing for QED or SA. (See details in Appendix C.1.2). We observe such a trade-off between QED and SA with Vina Dock in our ablation studies. We also experiment with various fine-tuning steps $n$. We show that Vina Dock metric will continue to improve with more fine-tuning steps.